Crew exchange, DV visit, and sediment cores – University of Copenhagen

Forward this page to a friend Resize Print Bookmark and Share

NEEM > Field diaries > Field season 2012 > Crew exchange, DV visi...

12 June 2012

Crew exchange, DV visit, and sediment cores

A group of interested DVs being guided around camp by the man in yellow also known as JP

Today we received a C-130 from Sønderstrømfjord (SFJ) that brought us new camp personnel plus a large group of interested Distinguished Visitors. After a two-hour tour of camp the DVs continued with departing camp members to Summit Camp for a stopover before returning to SFJ.

At the same time the deep drilling continued in the drill trench where several cores were retrieved. The core material is now a mixture of ice and ‘frozen mud’
apparently consisting of very fine grained material.

At the NEGIS site the shallow drilling continued successfully and reached 61m depth. Drillers were left alone at NEGIS while Polar 6 was shuttling NEGIS equipment.

What we have done today:
1. Drilling ‘ice-mud’ cores with rock drill in the deep hole
2. Receiving Skier 31 with DVs, NEEMers, cargo, and fuel
3. Providing DVs a 2-hour tour of camp
4. Two Polar 6 flights to NEGIS bringing equipment
5. Shallow drilling at NEGIS: 17m – 61m depth
6. Breaking down arriving pallets
7. NEGIS team working on seismic profile West of NEEM camp

Ad. 1: Subglacial drilling continues slowly but steady with two runs per day. The retrieved material now contains less ice and more fine grained clay-like material. Cores are continuously handled and packed in dark/red light in order to allow for luminescence dating.

Ad. 2: Skier 01 arrived at NEEM at 11:06 with 3 NEEMers, 14 DVs (members of the Danish parliament, Forsvarskommandoen, Rigsombudsmanden, Greenland police officers, CPS personnel) and two media persons (Euronews). We received two pallets with scientific gear and food, 2 skidoos, sledges, and 5000 lbs fuel from the skier. Skier departed for Summit Camp at 13:09 with all DVs and 5 NEEMers using only half the skiway. Thank you all visitors, air guards, and camp personnel for a smooth operation.

Ad.5: For unknown reasons the new power supply broke shortly after engaging the drill motor. Drilling was, however, successfully continued with a spare power supply and drillers had good progress. If weather permits, drilling will continue tomorrow.

Weather: Beautiful day with clear blue sky and a few fog banks in the morning.
Temp. –9°C to – 21°C, 5-10 knots from S.

FL, Anders Svensson


← Previous entry                                                                                Next entry →